Proof of Bose-Einstein Condensation for Dilute Trapped Gases

Robert Seiringer Princeton University

The ground state of bosonic atoms in a trap has been shown experimentally to display Bose-Einstein condensation (BEC). We prove this fact theoretically for bosons with two-body repulsive potentials in the dilute limit, starting from the basic Schrödinger equation; the condensation is 100that minimizes the Gross-Pitaevskii energy functional. This is the first rigorous proof of BEC in a physically realistic, continuum model.